Grade 1210 grade maths

A train travels 360 km at a uniform speed. If the speed had been 5 km/h more, it would have taken 1 hour less for the same journey. Find the speed of the train

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let
the speed of the train bexkm/hr.

Time taken to cover 360 km = 360/x

As per the question given,

⇒(x+ 5)(360-1/x) = 360

⇒360 -x+ 1800-5/x = 360

⇒x(x+ 45) -40(x+ 45) = 0

⇒(x+ 45)(x- 40) = 0

⇒x= 40, -45

As we know, the value of speed cannot be negative.

Therefore, the speed of train is 40 km/h.

Thanks